OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (-7, 6) and is perpendicular to -7x + 9y = -2

OpenStudy (anonymous):

hi you can change the equation to the form y=mx+b and your slope is m, b intercept

OpenStudy (anonymous):

-7x + 9y = -2 9y = -2 + 7x or write it as 9y= 7x-2 7x - 2 y= ---- --- 9 9 here your slope m= 7/9 right? but you prob wants a perpendicular to your line y=(7/9)x-(2/9) in order to be perpendicular ti this line your m2= -1/m1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so we get m2= -1/m1 m2= -1/(7/9) m2= -9/7 now that we know its slope we can form an equation of that line perpendicular to the first one

OpenStudy (anonymous):

okay am following your lead

OpenStudy (anonymous):

using the formula m=(y2-y1)/(x2-x1) or we can arrange them to y2-y1= m ( x2 - x1) just plug in your points(-7,6)=(x1,y1) here x1=-7 and y1=6 y2-6 =(-9/7)[x2-(-7)] now just change y2=y and x2=x y - 6 =(-9/7)[ x -(-7)] y - 6 =(-9/7)( x + 7 ) now do multiplication and subtraction y-6= -(9/7)x + (-9/7)7 y-6= -(9/7)x - 7 y = -(9/7)x - 7+6 y = -(9/7)x - 1 ans equation of the line perpendicular to the first line equation

OpenStudy (anonymous):

let me hear from you that you got it or not..lol

OpenStudy (anonymous):

sry i was checking another problem

OpenStudy (anonymous):

well i am looking at it and my mind is playing catch up